Evaluate the given expression and express the result using the usual format for writing numbers (instead of scientific notation) 35C3.
Answer by stanbon(75887)

Evaluate the given expression and express the result using the usual format for writing numbers (instead of scientific notation)
35C3 = (35!)/[(35-3)!*3!] = (35!)/[(32!*3!) = (35*34*33)/(1*2*3) = 6545
=====================================
Cheers,
Stan H.
=================